The average height of students in a freshman class of a certain school has been 149.45 cm with a population standard deviation of 7.74 cm. Is there a reason to believe that there has been a change in the average height if a random sample of 43 students in the present freshman class has an average height of 164 cm? Use a 0.01 level of significance.






1
Expert's answer
2022-06-03T13:08:01-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=149.45"

"H_1:\mu\not=149.45"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a two-tailed test is "z_c = 2.5758."

The rejection region for this two-tailed test is "R = \{z:|z|>2.5758\}."

The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{164-149.45}{7.74/\sqrt{43}}=12.3270"

Since it is observed that "|z|=12.3270>2.5758=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(z>12.3270)= 0," and since "p= 0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 4.25, at the "\alpha = 0.01" significance level.